OpenStudy (anonymous):

solve by the elimination method 3x-2y=0 -3x-y=-18

OpenStudy (mathteacher1729):

The second equation says y = -3x + 18 So input that into the first equation and solve for y. Once you have the value for y, put that into either equation and solve for x. :)

OpenStudy (anonymous):

it's easier to put it on top of each other so you can see the positive's and negatives 3x-2y=0 -3x-y=-18 Cancel out -3's, and so you would have -3y=-18. y=6. Substitute 6 into original equation. And mathteacher, he asked for elimination, not substitution.

OpenStudy (mathteacher1729):

Some1, thank you for the correction. I am on autopilot. Time to sign off for the evening. :) Sorry for the confusion Mr.Crazy.

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

Its ok thanks